\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{10}

\bibitem{AL07}
Alexandru Aleman, \emph{A class of integral operators on spaces of analytic
  functions}, Topics in complex analysis and operator theory, Univ. M\'{a}laga,
  M\'{a}laga, 2007, pp.~3--30.

\bibitem{AC01}
Alexandru Aleman and Joseph~A. Cima, \emph{An integral operator on {$H^p$} and
  {H}ardy's inequality}, J. Anal. Math. \textbf{85} (2001), 157--176.

\bibitem{AC09}
Alexandru Aleman and Olivia Constantin, \emph{Spectra of integration operators
  on weighted {B}ergman spaces}, J. Anal. Math. \textbf{109} (2009), 199--231.

\bibitem{APR19}
Alexandru Aleman, Sandra Pott, and Mar\'{\i}a~Carmen Reguera,
  \emph{Characterizations of a limiting class {$B_\infty$} of
  {B}\'{e}koll\'{e}-{B}onami weights}, Rev. Mat. Iberoam. \textbf{35} (2019),
  no.~6, 1677--1692.

\bibitem{Ale_Arist}
Alexandru Aleman and Aristomenis~G. Siskakis, \emph{An integral operator on
  {$H^p$}}, Complex Variables Theory Appl. \textbf{28} (1995), no.~2, 149--158.

\bibitem{Volberg98}
Krzysztof Barański, Alexander Volberg, and Anna Zdunik, \emph{Brennan's
  conjecture and the {M}andelbrot set}, International Mathematics Research
  Notices \textbf{1998} (1998), 589--600.

\bibitem{BM07}
A.~F. Beardon and D.~Minda, \emph{The hyperbolic metric and geometric function
  theory}, Quasiconformal mappings and their applications, Narosa, New Delhi,
  2007, pp.~9--56.

\bibitem{BE81}
David B\'{e}koll\'{e}, \emph{In\'{e}galit\'{e} \`a poids pour le projecteur de
  {B}ergman dans la boule unit\'{e} de {${\bf C}\sp{n}$}}, Studia Math.
  \textbf{71} (1981/82), no.~3, 305--323.

\bibitem{BER99}
Daniel Bertilsson, \emph{On {B}rennan's conjecture in conformal mapping}, Royal
  Institute of Technology, Stockholm, 1999.

\bibitem{betsakos2024monotonicity}
Dimitrios Betsakos and Nikolaos Karamanlis, \emph{On the monotonicity of the
  speeds for semigroups of holomorphic self-maps of the unit disk},
  Transactions of the American Mathematical Society \textbf{377} (2024),
  no.~02, 1299--1319.

\bibitem{BCDM20}
Filippo Bracci, Manuel~D. Contreras, and Santiago D\'{\i}az-Madrigal,
  \emph{Continuous semigroups of holomorphic self-maps of the unit disc},
  Springer Monographs in Mathematics, Springer, Cham, 2020.

\bibitem{bracci2023complete}
Filippo Bracci, Eva~A Gallardo-Guti{\'e}rrez, and Dmitry Yakubovich,
  \emph{Complete frequencies for koenigs domains}, arXiv:2311.17470 (2023).

\bibitem{bracci2020semigroup}
Filippo Bracci and Oliver Roth, \emph{Semigroup-fication of univalent self-maps
  of the unit disc}, arXiv:2002.08252 (2020).

\bibitem{BRE78}
James~E. Brennan, \emph{The integrability of the derivative in conformal
  mapping}, J. London Math. Soc. (2) \textbf{18} (1978), no.~2, 261--272.

\bibitem{carleson1994some}
Lennart Carleson and Nikolai~G Makarov, \emph{Some results connected with
  brennan's conjecture}, Arkiv f{\"o}r Matematik \textbf{32} (1994), no.~1,
  33--62.

\bibitem{nikos24}
Nikolaos Chalmoukis and Georgios Nikolaidis, \emph{On the boundedness of
  generalized integration operators on hardy spaces}, arXiv: 2405.13920 (2024).

\bibitem{CON10}
Olivia Constantin, \emph{Carleson embeddings and some classes of operators on
  weighted {B}ergman spaces}, J. Math. Anal. Appl. \textbf{365} (2010), no.~2,
  668--682.

\bibitem{contreras2024slope}
Manuel~D Contreras, Francisco~J Cruz-Zamorano, and Luis Rodr{\'\i}guez-Piazza,
  \emph{The slope problem in discrete iteration}, arXiv:2406.08389 (2024).

\bibitem{contreras2023centralizers}
Manuel~D Contreras, Santiago D{\'\i}az-Madrigal, and Pavel Gumenyuk,
  \emph{Centralizers of non-elliptic univalent self-maps and the embeddability
  problem in the unit disc}, arXiv:2311.04134 (2023).

\bibitem{contreras2023uniform}
Manuel~D Contreras, Carlos G{\'o}mez-Cabello, and Luis Rodr{\'\i}guez-Piazza,
  \emph{On the uniform convergence of continuous semigroups}, arXiv:2304.12759
  (2023).

\bibitem{contpel}
Manuel~D. Contreras, Jos\'{e}~A. Pel\'{a}ez, Christian Pommerenke, and Jouni
  R\"{a}tty\"{a}, \emph{Integral operators mapping into the space of bounded
  analytic functions}, J. Funct. Anal. \textbf{271} (2016), no.~10, 2899--2943.

\bibitem{FMCG76}
Jinfu Feng and Thomas~H. Macgregor, \emph{Estimates on integral means of the
  derivatives of univalent functions}, Journal d’Analyse Math{\'e}matique
  \textbf{29} (1976), 203--231.

\bibitem{CR85}
Jos\'{e} Garc\'{\i}a-Cuerva and Jos\'{e}~L. Rubio~de Francia, \emph{Weighted
  norm inequalities and related topics}, North-Holland Mathematics Studies,
  vol. 116, North-Holland Publishing Co., Amsterdam, 1985, Notas de
  Matem\'{a}tica, 104. [Mathematical Notes].

\bibitem{GirPel06}
Daniel Girela and Jos\'{e}~\'{A}ngel Pel\'{a}ez, \emph{Carleson measures,
  multipliers and integration operators for spaces of {D}irichlet type}, J.
  Funct. Anal. \textbf{241} (2006), no.~1, 334--358.

\bibitem{GOL69}
G.~M. Goluzin, \emph{Geometric theory of functions of a complex variable},
  Translations of Mathematical Monographs, vol. Vol. 26, American Mathematical
  Society, Providence, RI, 1969.

\bibitem{gumenyuk2024angular}
Pavel Gumenyuk, Maria Kourou, and Oliver Roth, \emph{The angular derivative
  problem for petals of one-parameter semigroups in the unit disk}, Revista
  Matem{\'a}tica Iberoamericana (2024).

\bibitem{HH2002}
H{\aa}kan Hedenmalm, \emph{The dual of a {B}ergman space on simply connected
  domains}, vol.~88, 2002, Dedicated to the memory of Tom Wolff, pp.~311--335.

\bibitem{HKZ00}
H{\aa}kan Hedenmalm, Boris Korenblum, and Kehe Zhu, \emph{Theory of {B}ergman
  spaces}, Graduate Texts in Mathematics, vol. 199, Springer-Verlag, New York,
  2000.

\bibitem{HS05}
H{\aa}kan Hedenmalm and Serguei Shimorin, \emph{Weighted {B}ergman spaces and
  the integral means spectrum of conformal mappings}, Duke Math. J.
  \textbf{127} (2005), no.~2, 341--393.

\bibitem{alan2}
H{\aa}kan Hedenmalm and Alan Sola, \emph{Spectral notions for conformal maps: a
  survey}, Comput. Methods Funct. Theory \textbf{8} (2008), no.~1-2, 447--474.

\bibitem{kourou2024rates}
Maria Kourou, Eleftherios~K Theodosiadis, and Konstantinos Zarvalis,
  \emph{Rates of convergence for holomorphic semigroups of finite shift},
  arXiv:2403.06883 (2024).

\bibitem{makarov1999fine}
Nikolai~G Makarov, \emph{Fine structure of harmonic measure}, ST PETERSBURG
  MATHEMATICAL JOURNAL C/C OF ALGEBRA I ANALIZ \textbf{10} (1999), 217--268.

\bibitem{Pel23}
Jos\'{e}~\'{A}ngel Pel\'{a}ez, Jouni R\"{a}tty\"{a}, and Fanglei Wu,
  \emph{Volterra-type operators mapping weighted {D}irichlet space into
  {$H^\infty$}}, Math. Z. \textbf{304} (2023), no.~2, Paper No. 31, 34.

\bibitem{POM92}
Ch. Pommerenke, \emph{Boundary behaviour of conformal maps}, Grundlehren der
  mathematischen Wissenschaften [Fundamental Principles of Mathematical
  Sciences], vol. 299, Springer-Verlag, Berlin, 1992.

\bibitem{POM75}
Christian Pommerenke, \emph{Univalent functions}, Studia
  Mathematica/Mathematische Lehrb\"{u}cher [Studia Mathematica/Mathematical
  Textbooks], vol. Band XXV, Vandenhoeck \& Ruprecht, G\"{o}ttingen, 1975, With
  a chapter on quadratic differentials by Gerd Jensen.

\bibitem{Rt07}
Jouni R\"{a}tty\"{a}, \emph{Integration operator acting on {H}ardy and weighted
  {B}ergman spaces}, Bull. Austral. Math. Soc. \textbf{75} (2007), no.~3,
  431--446.

\bibitem{SHAP93}
Joel~H. Shapiro, \emph{Composition operators and classical function theory},
  Universitext: Tracts in Mathematics, Springer-Verlag, New York, 1993.

\bibitem{SHI05}
Serguei Shimorin, \emph{Weighted composition operators associated with
  conformal mappings}, Quadrature domains and their applications, Oper. Theory
  Adv. Appl., vol. 156, Birkh\"{a}user, Basel, 2005, pp.~217--237.

\bibitem{Volberg17}
Wayne Smith, Dmitriy~M. Stolyarov, and Alexander Volberg, \emph{Uniform
  approximation of {B}loch functions and the boundedness of the integration
  operator on {$H^\infty$}}, Advances in Mathematics \textbf{314} (2017),
  185--202.

\bibitem{alan}
Alan Sola, \emph{An estimate of the universal means spectrum of conformal
  mappings}, Comput. Methods Funct. Theory \textbf{6} (2006), no.~2, 423--436.

\bibitem{zarvalis2023quasi}
Konstantinos Zarvalis, \emph{Quasi-geodesics and backward orbits under
  semigroups of holomorphic functions}, Illinois Journal of Mathematics
  \textbf{67} (2023), no.~1, 185--202.

\end{thebibliography}

\documentclass{amsart} 
\usepackage[utf8]{inputenc}
\usepackage{amsthm}
\usepackage{amsmath}

\usepackage{enumerate}   
\usepackage{amssymb}
\usepackage{fancyhdr} 
\usepackage{xcolor}
\newcommand{\norm}[1]{\left\lVert#1\right\rVert}
\usepackage{graphicx}
\usepackage{ esint }
\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{lemma}[theorem]{Lemma}
\theoremstyle{definition}

\theoremstyle{remark}

\theoremstyle{remark}
\newtheorem*{Littlewood}{Littlewood's subordination principle}
\theoremstyle{remark}
\newtheorem*{Koebe}{Koebe's growth and distortion theorem}
\renewcommand\Re{\operatorname{Re}}
\renewcommand\Im{\operatorname{Im}}

\usepackage{url}
\usepackage[hidelinks]{hyperref}
\hypersetup{
	colorlinks=true,
	linkcolor=cyan,
	filecolor=cyan,
	citecolor =cyan,      
	urlcolor=cyan,
}
\author{Alexandru Aleman}
\address{Lund University, Mathematics, Faculty of Science, P.O. Box 118, S-221 00 Lund, Sweden}
\email{alexandru.aleman@math.lu.se}
\author{Athanasios Kouroupis}
\address{Department of mathematics, KU Leuven, Celestijnenlaan 200B, 3001, Leuven, Belgium}
\email{athanasios.kouroupis@kuleuven.be}

\title{Brennan's conjecture holds for semigroups of holomorphic functions}
\date{}

\begin{document}

\begin{abstract}
In the present note, we give a short proof of Brennan's conjecture in the special case of continuous semigroups of holomorphic functions. We apply classical techniques of complex analysis in conjunction with recent results on B\'{e}koll\'{e}-Bonami weights and spectra of integration operators.
\end{abstract}
\maketitle
\section{Introduction}
J. E. Brennan in \cite{BRE78} conjectured that if $f:\mathbb{D}\mapsto \mathbb{C}$ is a conformal map. Then, for every $p\in(-2,\frac{2}{3})$ the $p$-integral mean of the derivative is finite:
\begin{equation}\label{eq:brennan}
\int\limits_{\mathbb{D}}\left|f'(w)\right|^p\,dA(w)<\infty,
\end{equation}
where $dA(z)=\frac{dx\,dxy}{\pi}$ is the normalized Lebesgue measure of the unit disk $\mathbb{D}$.

Brennan's conjecture is one of the most famous remaining open problems in the field of geometric function theory. The best result so far is due to Alan Sola \cite{alan2, alan}, where he proved that the conjecture holds for the values $p\in(-p_0,\frac{2}{3}),\,p_0\sim 1.76$. We refer the interested reader also to the article of Hedenmalm and Shimorin \cite{HS05}. For further examples of univalent functions satisfying the conjecture and connections with complex dynamics, we refer to \cite{Volberg98}.      

In the present note, we will restrict ourselves to conformal maps $f$ that can be embedded into a continuous semigroup of holomorphic functions in the unit disk.
\begin{theorem}\label{main}
Let $\{\phi_t(z)\}_{t\geq 0}$ be a continuous semigroup of conformal self-maps of the unit disk. Then, for every $p\in(-2,\frac{2}{3})$
\begin{equation}\label{eq:sembrennan}
\int\limits_{\mathbb{D}}\left|\phi_t'(w)\right|^p\,dA(w)<\infty.
\end{equation}
\end{theorem} 
A (continuous) semigroup of holomorphic functions, \cite{BCDM20}, is a family $\{\phi_t(z)\}_{t\geq 0}$ of conformal self-maps of the unit disk, such that:
\begin{enumerate}[(i)]
	\item $\phi_0(z)=z,\,z\in\mathbb{D}$.
	\item $\phi_{t+s}(z)=\phi_t\circ\phi_s(z),\,z\in\mathbb{D}$.
	\item For every $z\in\mathbb{D}$, the map $t\mapsto \phi_t(z)$ is continuous for $t\geq0$.
\end{enumerate}
%The result is optimal in the sense that the non-trivial iterates of the semigroup $\phi_t(z)= k^{-1}\left(k(z)+t\right)$, where $k(z)=\frac{z}{(1-z)^2}$ is the Koebe function, satisfy \eqref{eq:sembrennan} only for  $p\in(-2,\frac{2}{3})$.
\subsection*{Motivation}  Bertilsson, see \cite{BER99,HH2002}, proved that Brennan's conjecture is equivalent to the following estimate:
\[\int_0^{2\pi}|f'(re^{i\theta})|^{-2}\,d\theta\leq C\frac{|f'(0)|^{-2}}{1-r},\qquad f\text{ conformal},\qquad r\in(0,1),\]
where $C$ is an absolute constant. By a breakthrough result of N. Makarov the validity of Brennan's conjecture for bounded maps would imply the conjecture in its full generality, see \cite{carleson1994some,makarov1999fine}. Thus, for a potential proof of the conjecture it is sufficient to work only with slit-maps or their corresponding Löwner chains, \cite{GOL69}. But, what if we give a Löwner chain an extra additive structure?

This was exactly the reason why we are considering semigroups of holomorphic functions.
Note that a semigroup with Denjoy-Wolff point at the origin behaves like a Löwner chain, meaning that it  satisfies the differential equation
\[\frac{d\phi_t(z)}{dt}=-\phi_t(z)p(\phi_t(z)),\qquad\Re p(z)\geq 0,\]
and the following subordination relation
\[\phi_s\prec\phi_t,\qquad 0\leq t\leq s.\]

 The theory of semigroups in the unit disk has gained popularity the last years, \cite{betsakos2024monotonicity, bracci2023complete, contreras2024slope,contreras2023centralizers, contreras2023uniform, gumenyuk2024angular, kourou2024rates,zarvalis2023quasi}. It will be interesting to see if one can work towards Brennan's conjecture having as a starting point the result of this note and establishing semigroup-fication (\cite{bracci2020semigroup}) or approximation theorems between conformal functions and semigroups.

\subsection*{Acknowledgments}
The authors thank Manuel Contreras, Alan Sola and Georgios Stylogiannis for their helpful comments and discussions.

A. Kouroupis received support from Research Foundation--Flanders (FWO), Odysseus Grant No. G0DDD23N

\subsection*{Notation}
Throughout the paper, if $f,g: E\to \mathbb R$ are two functions defined on the same set $E$, the notation $f\lesssim  g$ will mean that there exists some constant $C > 0$ such that $f \leq Cg$ on $E$. We will write  $f \sim g$ when $f\lesssim  g$ and $g\lesssim  f$.
\section{Background Material}
\subsection{Holomorphic Semigroups and conformal functions}
We will make a short presentation on the topic and we refer the interested reader to \cite{BCDM20} for a thorough introduction and an overview of recent developments.

We recall that a holomorphic self-map of the unit disk $\phi$, which is not the identity, satisfies one of the following:
\begin{enumerate}[(i)]
\item $\phi$ has a unique fixed point $\tau$ in $\mathbb{D}$.\\
\item $\phi$ has no fixed point in $\mathbb{D}$ and there exists a unique point $\tau$ in the unit circle $\mathbb{T}$ such that
\[\angle\lim\limits_{z\rightarrow\tau}\phi(z)=\tau\qquad \text{and} \qquad \phi'(\tau):=\angle\lim\limits_{z\rightarrow\tau}\phi'(z)=\liminf_{z\rightarrow\tau}\frac{1-|\phi(z)|}{1-|z|}\leq 1.\]
\end{enumerate}
The point $\tau\in\overline{\mathbb{D}}$ is called the Denjoy-Wolff point of $\phi$. If $\tau\in\mathbb{D}$ then the map $\phi$ is called elliptic, if $\tau\in\mathbb{T}$ and $\phi'(\tau)=1$ it is called parabolic and finally if $\tau\in\mathbb{T}$ and $\phi'(\tau)<1$ it is called hyperbolic. 

If $\{\phi_t(z)\}_{t\geq 0}$ is a continuous non-trivial semigroup in $\mathbb{D}$ then all iterates different from the identity have the same Denjoy-Wolff point $\tau\in\overline{\mathbb{D}}$. Furthermore, if $\tau\in\mathbb{T}$ then
\[\phi_t'(\tau)=e^{-\lambda t},\qquad \Re \lambda\geq 0.\]
The semigroup $\{\phi_t(z)\}_{t\geq 0}$ is called elliptic, parabolic or hyperbolic if $\phi_1$ is elliptic, parabolic or hyperbolic, respectively. From now on we will assume that every semigroup is non-trivial and continuous.

For every semigroup $\{\phi_t(z)\}_{t\geq 0}$ of the unit disk there exists a conformal map $h(z)$ in the unit disk $\mathbb{D}$, the Koenings map, such that:
\begin{equation}
\phi_t(z)=h^{-1}\circ\psi_t\circ h(z),\qquad z\in \mathbb{D},
\end{equation}
where
\begin{enumerate}[(i)]
\item $\psi_t(z)= e^{-\lambda t}z,\, \Re\lambda\geq0$, when $\{\phi_t(z)\}_{t\geq 0}$ is an elliptic semigroup.\label{elliptic}
\item $\psi_t(z)= z+it,$ when $\{\phi_t(z)\}_{t\geq 0}$ is a non-elliptic semigroup (parabolic or hyperbolic).
\end{enumerate}
In the first case, the Koenig map $h$ is a $\lambda$-spirallike function with respect to $\tau$, that is 
\begin{equation}\label{koenigel}
\Re \left(\frac{1}{\lambda}(z-\tau)(1-\overline{\tau}z)\frac{h'(z)}{h(z)}\right)\geq 0.
\end{equation}

An equivalent geometric definition of a $\lambda$-spirallike function $h$ with respect to $\tau\in\mathbb{D}$ is the following:
\begin{enumerate}[(i)]
	\item The function $h$ is conformal in the unit disk and $h(\tau)=0$.
	\item If a point $z$ exists in the range $\Omega:=h(\mathbb{D})$ of the conformal map $h$, then the $\lambda$- spiral connecting $z$ and $0$ is a subset of $\Omega$.
	\[e^{-\lambda t}z\in\Omega,\qquad t\geq 0,\qquad z\in\Omega.\]

\end{enumerate}

If $\lambda\in(0,\infty)$ then the map $h$ is called starlike.

In the non-elliptic case, $h$ is starlike at infinity with respect to $\tau\in\mathbb{T}$:
\begin{equation}\label{koenignel}
\Im \left(\overline{\tau}(\tau-z)^2h'(z)\right)\geq 0,
\end{equation}
or equivalently
\begin{enumerate}[(i)]
	\item The function $h$ is conformal in the unit disk and $\limsup\limits_{z\rightarrow \tau}\Im h(z)=+\infty$.
	\item If a point $z$ exists in the range $\Omega:=h(\mathbb{D})$ of the conformal map $h$, then for every $t\geq 0$
	\[z+it\in\Omega.\]
	
\end{enumerate}
Our starting point was the observation that Brennan's conjecture holds for the Koening map of a semigroup. To prove this result, which is well known to the experts in this field, first we need to recall some classical theorems.
\begin{Littlewood}[\cite{SHAP93}]
Let $f$ be a holomorphic and $g$ be a conformal function in the unit disk with $f(\mathbb{D})\subset g(\mathbb{D})$ and $f(0)=g(0)$. Then for every $p>0$ and $r\in(0,1)$
\[\int\limits_0^{2\pi}|f(re^{i\theta})|^p\,d\theta\leq \int\limits_0^{2\pi}|g(re^{i\theta})|^p\,d\theta.\]
\end{Littlewood}

\begin{Koebe}[\cite{POM75}]
Let $f$ be a conformal map in the unit disk with $f(0)=0$ and $f'(0)=1$. Then
\begin{enumerate}[(i)]
\item \[\frac{|z|}{(1+|z|)^2}\leq |f(z)|\leq \frac{|z|}{(1-|z|)^2}.\]
\item \[\frac{1-|z|}{(1+|z|)^3}\leq |f'(z)|\leq \frac{1+|z|}{(1-|z|)^3}.\]
\item \[\frac{1-|z|}{1+|z|}\leq \frac{|zf'(z)|}{|f(z)|}\leq \frac{1+|z|}{1-|z|}.\]
\end{enumerate}
\end{Koebe}
A direct consequence of the Hardy-Stein identity is the following result for the integral means of a conformal map:
\begin{theorem}[\cite{POM92}]\label{integralmeans}
If $f$ is a conformal function in the unit disk and $f(z)=O\left((1-|z|)^{-a}\right),\, a\in (0,2]$ as $|z|\rightarrow 1^-$. Then for every $p>\frac{1}{a}$
\[\int\limits_0^{2\pi}|f(re^{i\theta})|^p\,d\theta= O\left(\frac{1}{(1-r)^{ap-1}}\right),\qquad r\rightarrow 1^-.\]
For $0<p<\frac{1}{a}$ the associated $p$-integral means are uniformly bounded.
\end{theorem}
A well-known result of this theory is that the conjecture holds for the classes of close-to-convex and starlike functions. For these results, we refer to the original paper by J. E. Brennan \cite{BRE78} in which they are attributed to B. Dahlberg and J. Lewis
\begin{proposition}\label{BrennanKoenig}
Let $\{\phi_t(z)\}_{t\geq 0}$ be a non-trivial continuous semigroup of conformal self-maps of the unit disk and $h$ be the associated Koenigs map. Then, for every $p\in(-2,\frac{2}{3})$
\begin{equation}
\int\limits_{\mathbb{D}}\left|h'(w)\right|^p\,dA(w)<\infty.
\end{equation}
\end{proposition}
\begin{proof}
Without loss of generality, we can assume that $p\in(-2,-1)$. By \eqref{koenigel}, \eqref{koenignel} and the Koebe distortion theorem there exist a holomorphic function $f$ with $\Re f\geq 0$ such that
\[|h'(z)|^p\lesssim  |f(z)|^{|p|},\qquad 0<\delta<|z|<1.\]
We consider the function $g(z)=\frac{\overline{f(0)}z+f(0)}{1-z}$ which maps the unit disk onto the right half-plane. We apply Littlewood's subordination principle and Theorem~\ref{integralmeans}, yielding to
\begin{align*}
\int\limits_{\mathbb{D}}\left|h'(w)\right|^p\,dA(w)&\lesssim  \int\limits_{\delta}^1\int\limits_{0}^{2\pi}\left|h'(re^{i\theta})\right|^p\,d\theta\,dr\\
&\lesssim \int\limits_{\delta}^1\int\limits_{0}^{2\pi}\left|g(re^{i\theta})\right|^{|p|}\,d\theta\,dr\\
&\lesssim \int\limits_{\delta}^1\frac{1}{(1-r)^{|p|-1}}\,dr<\infty.\qedhere
\end{align*}
\end{proof}

\subsection{B\'{e}koll\'{e}-Bonami weights and integration operators}
Given a non-negative, integrable function $\omega$ on the unit disk, we denote by $L^p(\omega),\,p>0$ the $L^p$ space with respect to the measure $\omega(z)\,dA(z)$. Let $L_\alpha^p(\omega)$ be the corresponding Bergman space of analytic functions, \cite{HKZ00}.

B\'{e}koll\'{e}-Bonami weights are the analogs of the well-known Muckenhoupt weights, \cite{CR85}.
For $p>1,\,\eta>-1$ the class of weights $B_p(\eta)$ consists of all weights $\omega$ such that
\begin{equation}\label{bekk}
\int\limits_{S(\theta,h)}\omega(z)\,dA_\eta(z)\left(\int\limits_{S(\theta,h)}\omega^\frac{-1}{p-1}(z)\,dA_\eta(z)\right)^{p-1}\lesssim  \left(A_\eta\left(S(\theta,h)\right)\right)^p,
\end{equation}
where $dA_\eta(z)=(1+\eta)\left(1-|z|^2\right)^\eta\,dA(z)$ and $S(\theta,h)$ is an arbitrary Carleson box
\begin{equation*}
S(\theta,h)=\left\{z=re^{it}: 1-h<r<1,\,\left|t-\theta\right|<\frac{h}{2} \right\}.
\end{equation*}
The implied constant in \eqref{bekk} is independent of the Carleson box.
If $\eta=0$ we will write $B_p$ instead of $B_p(0)$.

In \cite{BE81}, B\'{e}koll\'{e} proved that the Bergman projection 
\begin{equation*}
P_\eta(f)(z)=\int\limits_{\mathbb{D}}\frac{f(w)}{\left(1-\overline{z}w\right)^{\eta+2}}\,dA_\eta(z)\qquad \eta>-1,
\end{equation*}
is bounded from $L^p(\omega),\,p>1$ onto $L_\alpha^p(\omega)$ if and only if $\frac{\omega(z)}{\left(1-|z|^2\right)^\eta}\in B_p(\eta)$.

We list a couple of lemmata, essential for our purpose.
\begin{theorem}[\cite{AC09}]\label{AlCO}
Let $p_0>1,\,\eta>-1$ and $\omega(z)$ be a weight such that 
\[\frac{\omega(z)}{\left(1-|z|^2\right)^\eta}\in B_{p_0}(\eta).\]
Then, for every $p>0$ and every $f\in L_\alpha^p(\omega)$
\begin{equation}\label{LP2}
\int\limits_{\mathbb{D}} \left|f(z)\right|^p\omega(z)\,dA(z)\sim \left|f(0)\right|^p+\int\limits_{\mathbb{D}} \left|f'(z)\right|^p\left(1-|z|^2\right)^p\omega(z)\,dA(z).
\end{equation}
\end{theorem}
\begin{theorem}[\cite{CON10}]\label{CO}
Let $\mu$ be a positive Borel measure in the unit disk and $\omega(z)$ be a weight such that $\frac{\omega(z)}{\left(1-|z|^2\right)^\eta}\in B_{p_0}(\eta),\,p_0>1,\,\eta>-1$. Then, the property of $\mu$ being a Carleson measure for $L_\alpha^p(\omega)$ is independent of $p>0$.
\end{theorem}

A holomorphic self-map of the unit disk $\phi$ induces a (formal) composition operator $C_\phi(f)= f\circ \phi$. A direct consequence of the theorem above is that for B\'{e}koll\'{e}-Bonami weights $\omega(z)$, the property of $C_\phi$ to be bounded on $L_\alpha^p(\omega)$ is independent of $p>0$.
\subsection{Integration operators}
This is a well-studied topic that attracted the interest of plenty of mathematicians the last 30 years, see for example \cite{AL07,AC01, Ale_Arist,nikos24, contpel,GirPel06,Pel23,Rt07,Volberg17}.
Every function $g$ that belongs to the Bloch spaces corresponds to a bounded integration operator $T_g$ on $L_\alpha^p,\,p>0$:
\[T_g(f)=\int_0^zf(w)g'(w)\,dA(w).\]
Recall that the Bloch spaces $\mathfrak{B}$ contains all analytic functions $g:\mathbb{D}\mapsto \mathbb{C}$ that are Lipschitz continuous when the disk and the plane are equipped with the hyperbolic distance, see \cite{BM07}, and the euclidean metric, respectively. Or equivalently $\mathfrak{B}$ consists of all analytic functions $g$ in the unit disk such that
\[\sup_\mathbb{D}\{(1-|z|^2)|g'(z)|\}<\infty.\]
The spectrum of the operator $T_g$ has been determined in terms of B\'{e}koll\'{e}-Bonami weights in \cite{AC09,APR19}.
\begin{theorem}\label{tg}
 The point $\lambda\in\mathbb{C}\setminus\{0\}$ belongs to the resolvent set of $T_g$ acting on $L_\alpha^p,\, p\geq 1$ if and only if 
 \[e^{p\Re\frac{g}{\lambda}} \in B_\infty.\]
 If we further assume that the derivative of symbol $g$ is the Cauchy transform of a finite measure $\mu$ on the unit circle, that is:
\begin{equation}\label{cauchy}
g'(z)=\int_\mathbb{T}\frac{1}{\zeta-z}\,d\mu(\zeta),
\end{equation}
 then 
 \[\sigma\left(T_g\lvert_{L^p_\alpha}\right)=\{0\}\cup\overline{\left\{\lambda:e^{\frac{\Re g}{\lambda}}\notin L^p_\alpha\right\}}.\]
 
\end{theorem}
The class $B_\infty$ contains all weights $\omega$ in $\mathbb{D}$ such that
\[\sup\frac{1}{\int_{S(\theta,h)}\omega}\int_{S(\theta,h)}M\left(\omega\mathbf{1}_{S(\theta,h)}\right)<\infty,\]
where $M$ is the Hardy-Littlewood maximal function (over Carleson boxes) and the suppremum is taken over all Carleson boxes.

We extract the following lemma from \cite{APR19}:

\begin{lemma}\label{tglem}
 Let $g\in \mathfrak{B}$ and $r\in\mathbb{R}$. Then, the weight $\omega:=e^{r\Re g}$ belongs in $B_\infty$ if and only if there exists $q>1$ such that $\omega\in B_q$ if and only if there exists a $\gamma>0$ such that
 \begin{equation}
     \int_{\mathbb{D}}\frac{\omega(w)}{|1-\overline{z}w|^{\gamma+2}}\,dA(w)\lesssim  \frac{\omega(z)}{\left(1-|z|^2\right)^\gamma},\qquad z\in\mathbb{D}.
 \end{equation}
\end{lemma}

\section{Proof of the main result}
\begin{lemma}\label{bek}
The weight $\omega(z)=\left|h(z)\right|^{p},\,p\in(-2,-1)$, where $h$ is the Koenings map of a holomophic semigroup, exists in $B_q$ for some $q>1$.
\end{lemma}
\begin{proof}
Let us first assume that $h$ is starlike at infinity. For every automorphism of the unit disk $T_w(z)=\frac{w-z}{1-\overline{w}z},\,w\in\mathbb{D}$, there exists a holomorphic function $u:\mathbb{D}\rightarrow\{z:\Re z\geq0\}$ such that
\begin{equation*}
\left(h\circ T_w\right)'(z)=c_1\frac{u(z)}{(c_2-z)^2},\qquad c_1,\,c_2\in \mathbb{T},
\end{equation*}
one can prove that $c_1=iT_w(\tau)$ and $c_2=T_w(\tau)$, where $\tau$ is the Denjoy-Wolff point of the semigroup. By Littlewood's subordination principle applied as in Proposition~\ref{BrennanKoenig}, we have that
\begin{align*}
\int\limits_{\mathbb{D}} \left|\left(h\circ T_w\right)'(z)\right|^p\,dA(z)&\lesssim \int\limits_{\mathbb{D}}  \left|\frac{1}{u(z)}\right|^{|p|}\,dA(z)=\int\limits_{\mathbb{D}}  \left|\frac{\alpha-\overline{\alpha}z}{1-z}\right|^{|p|}\,dA(z)\\
&\leq C(p)|\alpha|^{|p|},
\end{align*}
where $|\alpha|=\left| \frac{1}{u(0)}\right|=  \frac{1}{\left|h'(w)\right|(1-|w|^2)}$.
Taking into account the behaviour of the automorphism $T_w$ in the Carleson box $S(\theta,h)$ with center $w=(1-\frac{h}{2})e^{i\theta}\in\mathbb{D}$, we obtain:
\begin{align*}
\frac{1}{\left|S(\theta,h)\right|}\int\limits_{S(\theta,h)}\left|h'(z)\right|^p\,dA(z)&\lesssim \left(1-|w|^2\right)^{|p|}\int\limits_{\mathbb{D}} \left|\left(h\circ T_w\right)'(z)\right|^p\,dA(z)\\
&\leq C(p)\left|h'(w)\right|^{p}.
\end{align*}
Working similarly or applying directly the following result of Feng and MacGregor \cite{FMCG76,HH2002}: For every conformal function $f$ in the unit disk and for every $b\in[0,\frac{2}{3})$
\[\int\limits_{\mathbb{D}}|f'(z)|^b\,dA(z)\lesssim  |f'(0)|^{b},\]
one can prove that
\begin{align*}
\frac{1}{\left|S(\theta,h)\right|}\int\limits_{S(\theta,h)}\left|h'(z)\right|^\frac{|p|}{3}\,dA(z)&\lesssim \left(1-|w|^2\right)^{\frac{p}{3}}\int\limits_{\mathbb{D}} \left|\left(h\circ T_w\right)'(z)\right|^{\frac{|p|}{3}}\,dA(z)\\
&\leq C(p)\left|h'(w)\right|^{\frac{|p|}{3}}.
\end{align*}
Thus, $|h'|^p\in B_4$.

For the spirallike case we can assume without loss of generality  that $h(0)=0$. By the Koebe distortion theorem the function $g(z)=\log\frac{h(z)}{z}$ belongs in the Bloch space. Furthermore, since $h$ is spirallike
\[h'(z)=\lambda \frac{h(z)}{z}u(z),\qquad \Re u\geq 0.\]
Herglotz representation formula \cite{POM92} yields to the existence of a finite measure $\mu$ on the unit circle such that
\[u(z)=\int_\mathbb{T}\frac{\zeta+z}{\zeta-z}d\mu(\zeta),\qquad z\in\mathbb{D},\]
and $\mu\left(\mathbb{D}\right)=\frac{1}{\lambda}$. Consequently, the derivative of the symbol $g$
is the Cauchy transform of a finite measure in the unit circle:
\[g'(z)=\frac{\lambda}{z}\left(u(z)-\frac{1}{\lambda}\right)=2\lambda\int_\mathbb{T}\frac{d\mu(\zeta)}{\zeta-z}.\]
The integration operator $T_g$ is bounded on $L^2_a$ and Theorem~\ref{tg} imply that $2\frac{r-p}{rp},$ where $r\in(-2,p)$, exists in the resolvent set of $T_g\lvert_{L^2_\alpha}$.

Therefore, by Lemma~\ref{tglem} $$\omega_1:=\left|\frac{h(z}{z}\right|^{\frac{pr}{r-p}}\in B_\infty,$$ and there exists a $\gamma>0$ such that

 \begin{equation}\label{part1}
     \int_{\mathbb{D}}\frac{\omega_1(w)}{|1-\overline{z}w|^{\gamma+2}}\,dA(w)\lesssim  \frac{\omega_1(z)}{\left(1-|z|^2\right)^\gamma},\qquad z\in\mathbb{D}.
 \end{equation}
Applying Littlewood's subordination principle in a similar manner us before, we obtain that 
 \begin{equation}\label{part2}
     \int_{\mathbb{D}}\frac{|u(w)|^r}{|1-\overline{z}w|^{4}}\,dA(w)\lesssim  \frac{|u
     (z)|^{r}}{\left(1-|z|^2\right)^2},\qquad z\in\mathbb{D}.
 \end{equation}
Let $\delta=2\frac{p}{r}+\gamma\frac{r-p}{r}>0$, Hölder's inequality yields to 
\begin{multline*}
     \int_{\mathbb{D}}\frac{\omega(w)(1-|z|^2)^{\delta}}{|1-\overline{z}w|^{\delta+2}}\,dA(w)\\\lesssim\left(\int_{\mathbb{D}}\frac{|u(w)|^r\left(1-|z|^2\right)^2}{|1-\overline{z}w|^{4}}\,dA(w)\right)^{\frac{p}{r}}\left(\int_{\mathbb{D}}\frac{\omega_1(w)\left(1-|z|^2\right)^\gamma}{|1-\overline{z}w|^{\gamma+2}}\,dA(w)\right)^{\frac{r-p}{r}} \\
  \lesssim \omega(z).
\end{multline*}

The proof follows applying again Lemma~\ref{tglem}, note that $\log h'\in\mathfrak{B}$.

\end{proof}
The two cases in the above lemma can be merged. The argument that we used when the Koenigs map is spirallike is far more general. Although, we decided to keep the first elementary argument for expository reasons.

The last ingredient that we need for the proof of Theorem~\ref{main} is an inequality similar to Schwarz-type lemmata for conformal functions. We will rely on a well-know technique of Sergei 
Shimorin, see for example \cite{SHI05} where he proved using the Löwner differential equation that every conformal self-map $\phi$ of the unit disk with $\phi(0)=0$, satisfies the following inequality:
\begin{equation}\label{shimorin}
\frac{1-|z|^2}{1-|\phi(z)|^2}\leq \left|\phi'(w)\right|\left(\frac{|z|}{|\phi(z)|}\right)^2.
\end{equation}

\begin{lemma}\label{ShimLow}
	Let $\phi:\mathbb{D}\mapsto\mathbb{D}$ be a conformal map that fixes the origin. Then, for every $p\in(0,2)$.
	\begin{equation}\label{lowner}
	\left|\phi'(z)\right|^p\geq \left(\frac{1-|z|^2}{1-|\phi(z)|^2}\right)^2\left|\frac{\phi(z)}{z}\right|^{2p}.
	\end{equation}
\end{lemma}
\begin{proof}
By a standard density argument, we can assume that $\phi$ is a slit mapping, meaning that $\phi(\mathbb{D})=\mathbb{D}\setminus\widehat{\gamma}$, where $\gamma$ is a smooth Jordan arc in $\mathbb{D}\setminus\{0\}$ with starting point in the unit circle.
There exists a Löwner chain (\cite{GOL69}) $\{\varphi_t\}_{0\leq t\leq T}$ such that
\begin{enumerate}[(i)]
\item $$\varphi_0(z)=z,\qquad \varphi_T(z)=\phi(z),\qquad z\in\mathbb{D}.$$
\item $$\frac{\partial}{\partial_t}\varphi_t(z)=-\varphi_t(z)\frac{1+k(t)\varphi_t(z)}{1-k(t)\varphi_t(z)},$$
where $k(t)$ is continuous with $|k(t)|=1$.
\end{enumerate}
We observe that:
\begin{align*}
\frac{\partial}{\partial_t}\log\frac{\left|\varphi'_t(z)\right|^{-p}\left|\frac{\varphi_t(z)}{z}\right|^{2p}}{\left(1-\left|\varphi_t(z)\right|^2\right)^2}&=-p\frac{\Re\left(\overline{\varphi'_t(z)}\frac{\partial}{\partial_t}\varphi'_t(z)\right)}{\left|\varphi'_t(z)\right|^2}\\
&+2p\frac{\Re\left(\overline{\varphi_t(z)}\frac{\partial}{\partial_t}\varphi_t(z)\right)}{\left|\varphi_t(z)\right|^2}+4\frac{\Re\left(\overline{\varphi_t(z)}\frac{\partial}{\partial_t}\varphi_t(z)\right)}{1-\left|\varphi_t(z)\right|^2}\\
&=\left(-p-4\frac{|\varphi_t(z)|^2}{1-|\varphi_t(z)|^2}\right)\Re \left(\frac{1+k(t)\varphi_t(z)}{1-k(t)\varphi_t(z)}\right)\\
&+p\Re\left(\frac{2k(t)\varphi_t(z)}{(1-k(t)\varphi_t(z))^2}\right)\\
&=-\left|1-k(t)\varphi_t(z)\right|^{-2}\left((4-p)\left|\varphi_t(z)\right|^2+p\right)\\
&+p\Re\left(\frac{2k(t)\varphi_t(z)}{(1-k(t)\varphi_t(z))^2}\right)\\
&\leq-\frac{(4-p)\left|\varphi_t(z)\right|^2-2p\left|\varphi_t(z)\right|+p}{\left|1-k(t)\varphi_t(z)\right|^{2}}\\
&\leq 0.
\end{align*}
Thus
\begin{equation*}
\frac{\left|\varphi'_t(z)\right|^{-p}\left|\frac{\varphi_t(z)}{z}\right|^{2p}}{\left(1-\left|\varphi_t(z)\right|^2\right)^2}\leq \frac{\left|\varphi'_0(z)\right|^{-p}\left|\frac{\varphi_0(z)}{z}\right|^{2p}}{\left(1-\left|\varphi_0(z)\right|^2\right)^2}=\left(1-|z|^2\right)^{-2}.
\end{equation*}
\end{proof}
Applying the above lemma in conjunction with the Koebe distortion theorem, we obtain:
\begin{lemma}\label{LOW}
For every conformal map $\phi:\mathbb{D}\mapsto\mathbb{D}$ and every $p\in(0,2)$, there exists a constant $C>0$ such that
\begin{equation}\label{brennaneq}
\left|\phi'(w)\right|^{\frac{p}{2}}\geq C \frac{1-|z|^2}{1-|\phi(z)|^2},\qquad z\in \mathbb{D}.
\end{equation}
\end{lemma}
\begin{proof}
Let us first assume that $\phi(0)=0$. Then, by Lemma~\ref{ShimLow} and Koebe's distortion theorem
\[\left|\phi'(z)\right|^{\frac{p}{2}}\geq \frac{1-|z|^2}{1-|\phi(z)|^2}\left|\frac{\phi(z)}{z}\right|^p\geq \left|\frac{|\phi'(0)|}{4}\right|^p\frac{1-|z|^2}{1-|\phi(z)|^2}.\]
If $\alpha=\phi(0)\neq0$ we consider the map $T_\alpha\circ\phi(z)=\frac{\alpha-\phi(z)}{1-\overline{\alpha}\phi(z)}$. By the previous case
 \[\left|\phi'(z)\right|^{\frac{p}{2}}\left|T_\alpha'\circ\phi(z)\right|^{\frac{p}{2}}\geq \left|\frac{|\phi'(0)|}{4(1-|\phi(0)|^2)}\right|^p\frac{1-|z|^2}{1-|\phi(z)|^2}\left|T_\alpha'\circ\phi(z)\right|^{-1}.\]
This completes the proof since $\left|T_\alpha'\circ\phi(z)\right|$ is bounded from above.
\end{proof}
\begin{proof}[\textbf{Proof of Theorem \ref{main}}]
We can assume that $\left\{\phi_t\right\}_{t\geq 0}$ is not a group and that $p\in(-2,-1)$.
\begin{align*}
\int\limits_{\mathbb{D}}\left|\phi_t'(w)\right|^p\,dA(w)&= C(t,p)\int\limits_{\mathbb{D}}\left|h'\circ\phi_t(w)\right|^{|p|}\left|h'(w)\right|^{p}\,dA(w)\\
&=C(t,p)\norm{C_{\phi_t}(h')}^{|p|}_{L^{|p|}(\omega)},
\end{align*}
where $\omega(z)=\left|h'(z)\right|^p$. By Lemma~\ref{bek} and Theorem~\ref{CO}, it is sufficient to prove that the composition operator $C_{\phi_t}$ is bounded on $L_\alpha^2(\omega)$. 

For an arbitrary function $f\in L_\alpha^2(\omega)$ we apply the Littlewood-Paley formula~\eqref{LP2}, yielding to:
\begin{align*}
&\norm{C_{\phi_t}(f)}_{L^2(\omega)}^2=\int\limits_{\mathbb{D}}\left|f\left(\phi_t\right)(w)\right|^2\left|h'(w)\right|^p\,dA(w)\\
&\lesssim  \left|f(\phi_t(0))\right|^2+\int\limits_{\mathbb{D}}\left|f'\left(\phi_t\right)(w)\right|^2\left|\phi_t'(w)\right|^2\left(1-|w|^2\right)^2\left|h'(w)\right|^p\,dA(w)\\
&\lesssim  \left|f(\phi_t(0))\right|^2+\int\limits_{\phi_t\left(\mathbb{D}\right)}\left|f'(w)\right|^2\left(1-|\phi_t^{-1}(w)|^2\right)^2\left|h'(\phi_t^{-1}(w))\right|^p\,dA(w).
\end{align*}
The point evaluation linear functionals are bounded since the weight $\omega$ is B\'{e}koll\'{e}-Bonami. Thus, applying again the Littlewood-Paley formula~\eqref{LP2} it is easy to see that the following inequality suffices boundedness of $C_{\phi_t}$ on $L_\alpha^2(\omega)$: 

\begin{equation*}
\left(1-|z|^2\right)^2\left|h'(z))\right|^p\lesssim \left(1-|\phi_t(z)|^2\right)^2\left|h'(\phi_t(z))\right|^p,\qquad z\in \mathbb{D},
\end{equation*}
or equivalently that
	\begin{equation}
\left|\phi_t'(z)\right|^{|p|}\geq C(\phi_t) \left(\frac{1-|z|^2}{1-|\phi_t(z)|^2}\right)^2.
\end{equation}
The proof follows from Lemma~\ref{LOW}
\end{proof}

	\bibliographystyle{amsplain-nodash} 
	\bibliography{refbre} 
\end{document}